find the set of three consecutive integers such that the first added to four times the third is 93
Answer by oberobic(2304)

3 consecutive integers can be defined as:
x
x+1
x+2
.
x + 4(x+2) = 93
.
x +4x +8 = 93
.
5x = 85
.
x = 17
.
Check for totaling 93.
.
17 + 4(17+2) = ?
17 + 4(19) = 93
Correct.
.
So the three consecutive integers are: 17, 18, and 19.
.
Done.